when did a charge become zero at an electric field having charges 16microcoulomb and -9mc.they have a distance of 8cm between them
 

Grade:12

1 Answers

Vikas TU
6 years ago
As we realize that 
 
E.F. = kq/r^2 
 
q = charge , r = separate 
 
Electric field will be zero somewhere close to them 
 
That is 
 
Net E.f. = k(16 x10^-6)/r^2 + k(9x10^-6)/(8-r)^2 
 
0 = k(16 x10^-6)/r^2 + k(9x10^-6)/(8-r)^2 
 
k(9x10^-6)/(8-r)^2 = k(16 x10^-6)/r^2 
 
=> 9/(8-r)^2 = 16/r^2 
 
=> 16(64+r^2-16r) = 9r^2 
 
From this we get,
 
=> r = 5.12